# Detecting Misuse of Security APIs: A Systematic Review

### v2

@@ 6.5.1. Heuristic-based
(iii) Hybrid Analysis: Attempts have been made to combine static and dynamic analysis in a hybrid approach to leverage the strengths of both techniques and overcome their weaknesses. Table 6 provides a concise summary of the strengths and weaknesses associated with static and dynamic analysis techniques. To mitigate the risk of false positives in the static analysis, some researchers proposed a hybrid approach that typically applies static analysis to identify potential misuses, followed by dynamic analysis to validate the results. Several studies [S8, S14, S19, S24, S57, S67] evaluated Android apps against a MitM attack to verify misuses reported by static analysis. Another study [S23] applied manual static analysis to find potential crypto misuses in Android apps and then performed dynamic memory analysis to examine the crypto libraries invoked during execution. This approach enables the detection of misuses that are feasible at runtime. Another study [S69] used a combination of static and dynamic analysis to find Android applications that call the SafetyNet Attestation API during their execution. Next, it did a manual static analysis to find vulnerable applications with potential misuses, followed by bypassing the SafetyNet Attestation checks to confirm the misuses.
Static analysis can also serve as a guide for dynamic analysis, reducing the time and memory consumption of dynamic analysis by pruning its exploration space. Some studies [e.g., S19, S24, S67] employed a preliminary static analysis to detect misuses. They further used static analysis for method call graphs to identify the entry points that trigger the execution of vulnerable methods. These entry points were then used to generate inputs for running applications during dynamic analysis, resulting in a more efficient analysis with a reduced input space. Another study [S6] combined static and dynamic analysis techniques to detect crypto misuses in iOS apps. It first used static analysis to find the locations of crypto APIs and then monitored those API calls at runtime using API hooking techniques. Misuses were detected by analyzing the execution logs, which record parameter values and other relevant information. AuthDroid [S47] also adopted a hybrid approach to detect OAuth misuse in Android apps. It uses static analysis to extract the basic elements of OAuth (e.g., user-agent, the identity of SP) from the app, then uses a MitM proxy in dynamic analysis to find API misuses in an authentication process. While the mentioned studies followed a static-dynamic approach in detecting misuses, one study [S26] has taken a different hybrid approach by first simulating a MitM attack to find vulnerable apps, and then manually performing static code analysis to identify the root causes of misuses.
@@ 6.5.2. ML-based
A. Feature Engineering: Three types of features were identified in the existing literature for building security API misuse detection models which are sequential, word-, and graph-based features.
@@
(iii) Graph-based features: One study [S65] utilized graph-based features to analyze the usage of security APIs in source code. First, the source code was parsed to AST and then modeled through graph embedding techniques, Bag of Graphs (BoG), and node2vec. These techniques are similar to word embedding techniques in NLP, where words are mapped to vectors based on how often they appear together in text. Similarly, BoG generates a collection of graph bag items representing elements or sub-graphs within a graph. These items are then used to construct a vector representation that captures the local attributes and relationships of the original graph. Node2vec is another graph embedding technique that extracts features from graphs, utilizing a flexible neighborhood sampling strategy.
B. Classification: In one study [S18], two probabilistic models, Hidden Markov Model (HMM) and n-gram, were trained using both secure and insecure API sequences. Notably, the labels for these sequences were generated using an existing tool, CogniCryptSAST [S11], designed for detecting misuses of cryptography APIs. The trained models were employed to predict the probability of a given API sequence being secure. An API sequence was considered insecure if its probability fell below a pre-defined threshold. The study also addressed the problem of identifying misuse locations within an insecure sequence by using a distance measure based on the probability of an API misuse at possible locations. In another study [S45], code snippets with the usage of security APIs were mined from SO and then classified using a Support Vector Machine (SVM) model. A small set of extracted code snippets was manually labeled to build the training dataset. Similarly, the approach proposed in [S65] used SVM, but trained a separate classifier for each misuse type, enabling the model to identify both the presence and type of API misuse. The classifiers were trained using labeled data containing correct and incorrect API use instances from two existing benchmarks for evaluating crypto misuse detection techniques, namely, Braga et al.’s (Braga et al., 2017) and Fischer et al.’s (Fischer et al., 2019) datasets.
@@ 7. RQ4: Evaluation methods
Our review identified 25 studies that performed experiments to assess the performance of misuse detection techniques. Out of these, 5 studies [S31, S32, S34, S41, S42] evaluated the performance of various static analysis tools in detecting crypto and SSL/TLS misuses. Following we discuss various metrics, benchmarks, and strategies adopted for evaluation.
@@ 7.1. Evaluation Strategies
Another evaluation technique, adopted by 3 studies, involves executing attacks to validate the results and identify exploitable misuses. For instance, one study [S6] executed two ethical attacks on two applications and successfully retrieved personal information encrypted and transmitted over the network. Meanwhile, 16 studies disclosed misuses identified in real-world projects, some of which analyzed the feedback they received from developers. This analysis provided valuable insights into developers’ requirements from misuse detection tools disregarded in existing approaches.
@@ 7.2. Evaluation Benchmarks
Benchmarks are critical for evaluating detection techniques and identifying their strengths and weaknesses. Unfortunately, there is an inadequate number of publicly available benchmarks, and all of them are limited to test cases for some misuses of Java crypto and SSL/TLS APIs. Table 7 lists 9 public benchmarks commonly used by researchers. Among them, the first five benchmarks were specifically created to evaluate and compare the performance of crypto misuse detection approaches.
CryptoAPI-Bench includes synthetic source codes with crypto API misuses, false positive tests, and correct API uses. It offers both basic test cases and advanced test cases that involve more complex scenarios. CryptoAPI-Bench was designed to assess static tools. CryptoAPI-Bench* [S10] is an extension of CrytoAPI-Bench with additional cases suitable for evaluating dynamic approaches. CryptoAPI-Bench is not suitable for evaluating the scalability property of a tool, as all test cases are lightweight by design. To address this limitation, another study [S32] created ApacheCryptoAPI-Bench using 10 real-world Apache projects that are complex programs with numerous and lengthy code files. This benchmark is therefore appropriate for assessing the scalability and applicability of existing approaches to real-world applications.
@@
R Benchmark apps for evaluating the performance of static information-flow analysis of Android apps including crypto misuse test cases S44
@@
R Benchmark apps for evaluating the performance of static analysis to detect inter-component data leakage problem of Android apps including crypto misuse test cases S44
@@ 7.3. Evaluation Metrics
Several primary studies also considered the computation efficiency of detection techniques, which was measured using the runtime and space complexity required for misuse detection. Evaluating computation efficiency is crucial in demonstrating the suitability of these techniques for real-world applications. Classification accuracy was used only by ML-based detection techniques, and coverage measurement is exclusive to dynamic analysis as it measures the proportion of the program code that has been executed during testing. Figure 13 illustrates the distribution of the identified evaluation metrics that are categorized by detection technique.
@@ 8.1.4. Lack of cross-language tool support
Along with the need to develop tools for various languages, there is a need for developing cross-language tools. A study [S20] analyzing MicroPython projects suggests that developers working in the embedded domain often use C language to implement cryptography operations. This indicates the need for tools with cross-language analysis capability to track program information across multiple programming languages. Furthermore, Meng et al.’s analysis of SO posts (Meng et al., 2018) has shed light on the challenge of cross-language data handling of cryptography APIs, where developers struggle to encrypt data in one language (e.g., PHP or Python) and decrypt data in another language (e.g., Java). To address these issues, future research should address the development of cross-language approaches to ensure secure software development across different programming languages.
@@ 8.2.1. Limited research on the application of ML-based techniques
Our SLR identified a critical gap: a scarcity of state-of-the-art ML-based models for identifying security API misuse. Over 95% of the reviewed studies rely on traditional heuristic-based approaches that require significant domain knowledge and result in labor-intensive, time-consuming, and error-prone processes. Manual analysis of identified misuses in one study [S58], using heuristic-based CrySL [S11], also revealed several false positives due to incorrect specifications defined in CrySL. Additionally, these methods often rely on hard-coded rules, making them difficult to adapt to emerging misuses to keep pace with the rapidly evolving security threat landscape. The main obstacle to adopting ML models in this domain is the scarcity of large-scale, labeled datasets. While some studies have manually analyzed and verified identified misuses, none have publicly shared the results of their analysis. We strongly encourage researchers to openly share their findings on security API usage analysis. This would not only support validation and future research, but would also significantly benefit the development of large-scale labeled datasets -crucial for both training ML models and establishing evaluation benchmarks.
@@ 8.2.3. Challenges in adapting models to API evolution
The ever-changing nature of APIs presents a significant challenge for misuse detection models. As APIs adapt to new requirements and security concerns, detection models must constantly keep pace to identify new misuse patterns arising from these changes. However, the rapid evolution often outpaces existing models’ ability to detect new forms of misuse. In our analysis, we noted the deprecation of two APIs, Fingerprint and SafetyNet Attestation, yet no solutions have addressed adapting misuse detection models to their alternatives, Biometrics (Google, a) and Play Integrity (Google, b) APIs. Another hurdle is the overreliance on hard-coded rules (over 95% of studies), which are inherently limited in detecting emerging misuse patterns. To overcome this, we recommend researchers adopt ML models and explore efficient incremental learning algorithms to address concept drift issues.
@@ 8.2.4. Lack of a standardized evaluation framework
Pendlebury et al. (Pendlebury et al., 2019) identified two critical sources of bias that inflate experimental results in ML-based malware classification: spatial bias and temporal bias. Spatial bias stems from unrealistic assumptions about the goodware-to-malware ratio within training and testing data. Temporal bias occurs due to unrealistic time splits within datasets, where future knowledge is included in the training data. Their proposed evaluation framework addresses these issues by incorporating spatial and temporal constraints. This concern extends beyond malware classification to any machine learning model in the security domain that deals with constantly evolving threats and time-sensitive data, including security API misuse detection. However, no reviewed ML-based studies considered experimental biases. To ensure accurate assessment of ML-based misuse detection in realistic settings, the research community must prioritize integrating spatial and temporal constraints into experimental design. This specifically requires augmenting datasets with accurate timestamps and acquiring knowledge of realistic ratios between normal and misuse classes.
@@ 8.3.3. Lack of support for repair
Developers may face challenges while trying to fix identified misuses, potentially introducing new mistakes [S42]. An analysis of vulnerability disclosures, reported by the reviewed studies, reveals instances where developers acknowledged misuses but struggled to resolve them due to several limitations. Operational constraints, like maintaining backward compatibility, can restrict their ability to implement necessary fixes [S1]. Additionally, existing tools often provide inadequate repair guidance, lacking the detailed information required for repair [S31]. Further, the inherent complexity of implementing secure solutions poses a significant challenge for developers [S31]. To address these challenges, misuse detection tools should be complemented with comprehensible, detailed, and actionable fixing suggestions. While some studies offer general repair guidance for specific misuse types [S5, S59, S63], they lack customizing fixes for a given vulnerable program. There are only four studies [S2, S16, S17, S46] for the automated generation of customized fixes for crypto misuses. Existing tools, however, are still inadequate in assisting developers with accurately correcting misuses [S31]. Future research should explore methods for providing detailed and customized suggestions and streamlining automated repair solutions.
@@ 9. Threats to Validity
We followed the guidelines outlined in study (Kitchenham and Charters, 2007) to design and conduct our SLR. However, there might still be some biases due to the author’s expertise and different perspectives. We took necessary steps to minimize the impact of potential biases including those in the study selection and review process, which are elaborated upon below:
@@
Data Extraction and Synthesize: Human errors and author biases during data extraction, analysis, and interpretation can impact the accuracy of the results and findings. To mitigate this issue, a data extraction form was developed and refined to ensure the collection of adequate and consistent information for answering the research questions. We also conducted fortnightly meetings among all the authors to review and verify the synthesis and interpretation of our quantitative and qualitative analysis. Any disagreements were resolved through this finalizing our responses to the RQs.
@@ 10. Conclusion
    We identified a total of 39 misuses, including 6 crypto, 6 SSL/TLs, 10 OAuth, 4 Fingerprint, 4 Spring, and 9 SafetyNet Attestation misuses. The primary studies mainly focused on analyzing Android apps, and the most commonly reported misuses were using insecure crypto algorithms for crypto APIs, improper certificate validation for SSL APIs, misuse of the state parameter for OAuth APIs, and lack of cryptography for Fingerprint APIs.
@@
Overall, our SLR offers valuable insights for both researchers and practitioners. Researchers can deepen their understanding of existing work and identify areas for their future studies. Practitioners benefit in two ways - firstly, by using the findings to develop more effective tools for detecting security API misuses and secondly, by gaining insights into potential misuses, enabling them to avoid them in their development processes. Ultimately, this review contributes to the broader goal of promoting secure software development.

### v3

@@ 6.5.1. Heuristic-based
(iii) Hybrid Analysis: Attempts have been made to combine static and dynamic analyses in a hybrid approach to leverage the strengths of both techniques and overcome their weaknesses. Table 6 provides a concise summary of the strengths and weaknesses associated with static and dynamic analysis techniques. To mitigate the risk of false positives in the static analysis, some researchers proposed a hybrid approach that typically applies static analysis to identify potential misuses, followed by dynamic analysis to validate the results. Several studies [S8, S14, S19, S24, S57, S67] evaluated Android apps against a MitM attack to verify misuses reported by static analysis. Another study [S23] applied manual static analysis to find potential crypto misuses in Android apps and then performed dynamic memory analysis to examine the crypto libraries invoked during execution. This approach enables the detection of misuses that are feasible at runtime. Another study [S69] used a combination of static and dynamic analyses to find Android applications that call the SafetyNet Attestation API during their execution. Next, it did a manual static analysis to find vulnerable applications with potential misuses, followed by bypassing the SafetyNet Attestation checks to confirm the misuses.
Static analysis can also serve as a guide for dynamic analysis, reducing the time and memory consumption of dynamic analysis by pruning its exploration space. Some studies [e.g., S19, S24, S67] employed a preliminary static analysis to detect misuses. They further used static analysis for method call graphs to identify the entry points that trigger the execution of vulnerable methods. These entry points were then used to generate inputs for running applications during dynamic analysis, resulting in a more efficient analysis with a reduced input space. Another study [S6] combined static and dynamic analysis techniques to detect crypto misuses in iOS apps. It first used static analysis to find the locations of crypto APIs and then monitored those API calls at runtime using API hooking techniques. Misuses were detected by analyzing the execution logs, which record parameter values and other relevant information. AuthDroid [S47] also adopted a hybrid approach to detect OAuth misuse in Android apps. It uses static analysis to extract the basic elements of OAuth (e.g., user-agent, the identity of SP) from the app, then uses a MitM proxy in dynamic analysis to find API misuses in the authentication process. While the mentioned studies followed a static-dynamic approach in detecting misuses, one study [S26] has taken a different hybrid approach by first simulating a MitM attack to find vulnerable apps, and then manually performing static code analysis to identify the root causes of misuses.
@@ 6.5.2. ML-based
A. Feature Engineering: Three types of features were identified in the existing literature for building security API misuse detection models, which are sequential, word-, and graph-based features.
@@
(iii) Graph-based features: One study [S65] utilized graph-based features to analyze the usage of security APIs in source code. First, the source code was parsed to an AST and then modeled through graph embedding techniques, Bag of Graphs (BoG), and node2vec. These techniques are similar to word embedding techniques in NLP, where words are mapped to vectors based on how often they appear together in text. Similarly, BoG generates a collection of graph bag items representing elements or sub-graphs within a graph. These items are then used to construct a vector representation that captures the local attributes and relationships of the original graph. Node2vec is another graph embedding technique that extracts features from graphs, utilizing a flexible neighborhood sampling strategy.
B. Classification: In one study [S18], two probabilistic models, Hidden Markov Model (HMM) and n-gram, were trained using both secure and insecure API sequences. Notably, the labels for these sequences were generated using an existing tool, CogniCryptSAST [S11], designed for detecting misuses of cryptography APIs. The trained models were employed to predict the probability of a given API sequence being secure. An API sequence was considered insecure if its probability fell below a pre-defined threshold. The study also addressed the problem of identifying misuse locations within an insecure sequence by using a distance measure based on the probability of an API misuse at possible locations. In another study [S45], code snippets with the usage of security APIs were mined from SO and then classified using a Support Vector Machine (SVM) model. A small set of extracted code snippets was manually labeled to build the training dataset. Similarly, the approach proposed in [S65] used SVM, but trained a separate classifier for each misuse type, enabling the model to identify both the presence and type of API misuse. The classifiers were trained using labeled data containing correct and incorrect API use instances from two existing benchmarks for evaluating crypto misuse detection techniques, namely datasets by Braga et al.  (Braga et al., 2017) and Fischer et al. (Fischer et al., 2019).
@@ 7. RQ4: Evaluation methods
Our review identified 25 studies that performed experiments to assess the performance of misuse detection techniques. Out of these, 5 studies [S31, S32, S34, S41, S42] evaluated the performance of various static analysis tools in detecting crypto and SSL/TLS misuses. Following, we discuss various metrics, benchmarks, and strategies adopted for evaluation.
@@ 7.1. Evaluation Strategies
Another evaluation technique, adopted by 3 studies, involves executing attacks to validate the results and identify exploitable misuses. For instance, one study [S6] executed two ethical attacks on two applications and successfully retrieved personal information encrypted and transmitted over the network. Meanwhile, 16 studies disclosed misuses identified in real-world projects, some of which analyzed the feedback they received from developers. This analysis provided valuable insights into developers’ requirements for misuse detection tools, disregarded in existing approaches.
@@ 7.2. Evaluation Benchmarks
Benchmarks are critical for evaluating detection techniques and identifying their strengths and weaknesses. Unfortunately, there is an inadequate number of publicly available benchmarks, and all of them are limited to test cases for some misuses of Java crypto and SSL/TLS APIs. Table 7 lists 9 public benchmarks commonly used by researchers. Among them, the first five benchmarks were specifically created to evaluate and compare the performance of misuse detection approaches for security APIs.
CryptoAPI-Bench includes synthetic source code examples with crypto API misuses, false positive tests, and correct API uses. It offers both basic test cases and advanced test cases that involve more complex scenarios. CryptoAPI-Bench was designed to assess static tools. CryptoAPI-Bench* [S10] is an extension of CryptoAPI-Bench with additional cases suitable for evaluating dynamic approaches. CryptoAPI-Bench is not suitable for evaluating the scalability of a tool, as all test cases are lightweight by design. To address this limitation, another study [S32] created ApacheCryptoAPI-Bench using 10 real-world Apache projects that are complex programs with numerous and lengthy code files. This benchmark is therefore appropriate for assessing the scalability and applicability of existing approaches to real-world applications.
@@
R Benchmark apps for evaluating the performance of static information-flow analysis of Android apps, including crypto misuse test cases S44
@@
R Benchmark apps for evaluating the performance of static analysis to detect inter-component data leakage problem of Android apps, including crypto misuse test cases S44
@@ 7.3. Evaluation Metrics
Several primary studies also considered the computation efficiency of detection techniques, which was measured using the runtime and space complexity required for misuse detection. Evaluating computation efficiency is crucial in demonstrating the suitability of these techniques for real-world applications. Classification accuracy was used only by ML-based detection techniques, and coverage measurement is exclusive to dynamic analysis, as it measures the proportion of the program code that has been executed during testing. Figure 13 illustrates the distribution of the identified evaluation metrics that are categorized by detection technique.
@@ 8.1.4. Lack of cross-language tool support
Along with the need to develop tools for various languages, there is a need to develop cross-language tools. A study [S20] analyzing MicroPython projects suggests that developers working in the embedded domain often use C language to implement cryptography operations. This indicates the need for tools with cross-language analysis capability to track program information across multiple programming languages. Furthermore, Meng et al.’s analysis of SO posts (Meng et al., 2018) has shed light on the challenge of cross-language data handling of cryptography APIs, where developers struggle to encrypt data in one language (e.g., PHP or Python) and decrypt data in another language (e.g., Java). To address these issues, future research should address the development of cross-language approaches to ensure secure software development across different programming languages.
@@ 8.2.1. Limited research on the application of ML-based techniques
Our SLR identified a critical gap: a scarcity of state-of-the-art ML-based models for identifying security API misuse. Over 95% of the reviewed studies rely on traditional heuristic-based approaches that require significant domain knowledge and result in labor-intensive, time-consuming, and error-prone processes. Manual analysis of identified misuses in one study [S58], using heuristic-based CrySL [S11], also revealed several false positives due to incorrect specifications defined in CrySL. Additionally, these methods often rely on hard-coded rules, making them difficult to adapt to emerging misuses to keep pace with the rapidly evolving security threat landscape. The main obstacle to adopting ML models in this domain is the scarcity of large-scale, labeled datasets. While some studies have manually analyzed and verified identified misuses, none have publicly shared the results of their analysis. We strongly encourage researchers to openly share their findings on security API usage analysis. This would not only support validation and future research but would also significantly benefit the development of large-scale labeled datasets, crucial for both training ML models and establishing evaluation benchmarks.
@@ 8.2.3. Challenges in adapting models to API evolution
The ever-changing nature of APIs presents a significant challenge for misuse detection models. As APIs adapt to new requirements and security concerns, detection models must constantly keep pace to identify new misuse patterns arising from these changes. However, the rapid evolution often outpaces existing models’ ability to detect new forms of misuse. In our analysis, we noted the deprecation of two APIs, Fingerprint and SafetyNet Attestation, yet no solutions have addressed adapting misuse detection models to their alternatives, Biometrics (Google, a) and Play Integrity (Google, b) APIs. Zhong and Meng’s recent work (Zhong and Meng, 2024) takes a step toward addressing this issue by introducing a compiler-directed approach that automatically updates callsites when compilation errors indicate that a previously valid API usage has become incompatible. This offers a promising direction for bridging the gap between evolving APIs and misuse detection models when changes result in compilation failures. However, many evolution scenarios—particularly in the security domain—do not manifest as compilation errors. Future research should therefore extend automated migration techniques to handle such cases and integrate these techniques with misuse detection models to ensure their continued effectiveness as APIs evolve. Another significant issue is the overreliance on hard-coded rules (over 95% of studies), which limits the ability to detect emerging misuse patterns. We therefore recommend that future work investigate ML-based models and efficient incremental learning algorithms to address concept drift as APIs continue to evolve.
@@ 8.2.4. Lack of a standardized evaluation framework
Pendlebury et al. (Pendlebury et al., 2019) identified two critical sources of bias that inflate experimental results in ML-based malware classification: spatial bias and temporal bias. Spatial bias stems from unrealistic assumptions about the goodware-to-malware ratio within training and testing data. Temporal bias occurs due to unrealistic time splits within datasets, where future knowledge is included in the training data. Their proposed evaluation framework addresses these issues by incorporating spatial and temporal constraints. This concern extends beyond malware classification to any machine learning model in the security domain that deals with constantly evolving threats and time-sensitive data, including ML-based security API misuse detection. Furthermore, heuristic-based approaches should also be tested under conditions that reflect actual usage scenarios and the evolution of APIs over time. However, no reviewed studies explicitly considered such biases. To ensure accurate assessment of misuse detection in realistic settings, the research community must prioritize integrating spatial and temporal constraints into experimental design. This specifically requires augmenting datasets with accurate timestamps and obtaining realistic ratios between normal and misuse cases.
@@ 8.3.3. Lack of support for repair
Developers may face challenges while trying to fix identified misuses, potentially introducing new mistakes [S42]. An analysis of vulnerability disclosures, reported by the reviewed studies, reveals instances where developers acknowledged misuses but struggled to resolve them due to several limitations. Operational constraints, like maintaining backward compatibility, can restrict their ability to implement necessary fixes [S1]. Additionally, existing tools often provide inadequate repair guidance, lacking the detailed information required for repair [S31]. Further, the inherent complexity of implementing secure solutions poses a significant challenge for developers [S31]. To address these challenges, misuse detection tools should be complemented with comprehensible, detailed, and actionable fixing suggestions. While some studies offer general repair guidance for specific misuse types [S5, S59, S63], they lack customized fixes for a given vulnerable program. There are only four studies [S2, S16, S17, S46] for the automated generation of customized fixes for crypto misuses. Existing tools, however, are still inadequate in assisting developers with accurately correcting misuses [S31]. Future research should explore methods for providing detailed and customized suggestions and streamlining automated repair solutions.
@@ 9. Threats to Validity
We followed the guidelines outlined in study (Kitchenham and Charters, 2007) to design and conduct our SLR. However, there might still be some biases due to the author’s expertise and different perspectives. We took necessary steps to minimize the impact of potential biases, including those in the study selection and review process, which are elaborated upon below:
@@
Data Extraction and Synthesis: Human errors and author biases during data extraction, analysis, and interpretation can impact the accuracy of the results and findings. To mitigate this issue, a data extraction form was developed and refined to ensure the collection of adequate and consistent information for answering the research questions. We also conducted fortnightly meetings among all the authors to review and verify the synthesis and interpretation of our quantitative and qualitative analysis. Any disagreements were discussed and resolved collaboratively before finalizing our responses to the RQs.
@@ 10. Conclusion
    We identified a total of 30 distinct types of security API misuses. The primary studies mainly focused on analyzing Android apps, and the most commonly reported misuses were using insecure crypto algorithms for crypto APIs, improper certificate validation for SSL APIs, lack or misuse of the state parameter for OAuth APIs, and lack or misuse of cryptography for Fingerprint APIs.
@@
Overall, our SLR offers valuable insights for both researchers and practitioners. Researchers can deepen their understanding of existing work and identify areas for their future studies. Practitioners benefit in two ways - firstly, by using the findings to develop more effective tools for detecting security API misuse and secondly, by gaining insights into potential misuses, enabling them to avoid them in their development processes. Ultimately, this review contributes to the broader goal of promoting secure software development.